OpenStudy (anonymous):

Sami consulted a dietitian regarding weight loss and was given a nutrition plan to follow. The dietitian informed Sami, if she follows the plan, her results should be in alignment with the following linear model. In this model, x represents the number of weeks on the plan, and y represents Sami's weight (in pounds). Interpret the y-intercept

OpenStudy (anonymous):

\[y=342-2.2x\]

OpenStudy (anonymous):

@Michele_Laino this is the equation it gave me

OpenStudy (anonymous):

A.The y-intercept represents Sami's goal weight upon completion of the plan. B.The y-intercept represents Sami's weight upon starting the plan. C.The y-intercept represents the amount of weight Sami should lose in total D.The y-intercept represents the amount of weight Sami should lose each week

OpenStudy (anonymous):

i guess what im asking is will u walk me through how to slove the equation

OpenStudy (michele_laino):

the y-intercept, is the weight y at x=0, so we have: y-intercept = 342-2.2*0= 342 Namely the y-intercept is the weight at the starting of the plan

OpenStudy (michele_laino):



OpenStudy (anonymous):

thats awsome but i dont understand

OpenStudy (michele_laino):

since the quantity x measures the number of weeks of the plan, then we have that at x=0, namely at the week number zero, the weight is: y=342- 2.2* 0= 342-0= 342

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

so that doesnt help me to solve the equation

OpenStudy (anonymous):

ugh i hate math and i suck at it to

OpenStudy (michele_laino):

please look at the second option

OpenStudy (anonymous):

k thanks

OpenStudy (michele_laino):

:)